The “lead” in Ballpen is a graphite composition with Young’s modulus of approximately 1.2×109 N/m2. Calculate the change in length of the lead in an automatic ballpen if you tap it straight into the ballpen with a force of 7.8 N. The lead is 0.55 mm in diameter and 80 mm long


1
Expert's answer
2021-03-25T14:08:58-0400

The change in length is


"\Delta l=\epsilon l."

The stress:


"\sigma=\frac FA=\frac{4F}{\pi d^2}."

On the other hand, this stress can be expressed as


"\epsilon=\frac \sigma E=\frac{4F}{\pi d^2E}."

Finally, the change in length is


"\Delta l=\epsilon l=\frac{4Fl}{\pi d^2E}=2.19\text{ mm}."